OpenStudy (anonymous):

Find the value of x. EF = 9x + 14, FG = 56, and EG = 250. I keep on getting the wrong answer..

OpenStudy (mathstudent55):

EG = 9x + 14 EG = 250 So 9x + 14 = 250 Solve for x Subtract 14 from both sides: 9x = 236 Now divide both sides by 9 x = 236/9

OpenStudy (anonymous):

I did that as well. However, the options they give are the following: x = 22 x = 20 x = 110 x = 180

OpenStudy (tamtoan):

what is FG stand for ?

OpenStudy (anonymous):

FG = 56

OpenStudy (tamtoan):

if it's just a number ..then they don't call FG ..something like ...h is the height in a triangle...

OpenStudy (mathstudent55):

Is there a picture that goes with this question? Is this a geometry problem?

OpenStudy (anonymous):

 Drawing is not to scale. F is not supposed to be in the middle of the line segment.

OpenStudy (anonymous):

I got it! :) x = 20

OpenStudy (anonymous):

9x + 14 + 56 = 250 9x + 70 = 250 Subtract 70 on both sides 9x = 180 Divide both sides by 9 x = 20

OpenStudy (tamtoan):

from the given information you type up there, it doesn't make any sense to me..

OpenStudy (anonymous):

Oh I meant to type that EF = 9x + 14. There was a typo I didn't notice. Does that clear up the confusion?

OpenStudy (tamtoan):

there :) now it makes sense

OpenStudy (anonymous):

Good. ^^ Thank you both for your help! You made me think more about the problem.